Lazorwolf64:

Help me pls 2(9) = 3e

Shalore100:

ombine like terms: 3e + -3e = 0 18 + -3e = 0 Add '-18' to each side of the equation. 18 + -18 + -3e = 0 + -18 Combine like terms: 18 + -18 = 0 0 + -3e = 0 + -18 -3e = 0 + -18 Combine like terms: 0 + -18 = -18 -3e = -18 Divide each side by '-3'. e = 6 Simplifying e = 6ombine like terms: 3e + -3e = 0 18 + -3e = 0 Add '-18' to each side of the equation. 18 + -18 + -3e = 0 + -18 Combine like terms: 18 + -18 = 0 0 + -3e = 0 + -18 -3e = 0 + -18 Combine like terms: 0 + -18 = -18 -3e = -18 Divide each side by '-3'. e = 6 Simplifying e = 6

Shalore100:

sorry it tookk too long i had too type

Shalore100:

what grade are you in

Lazorwolf64:

8th what about you?

Lazorwolf64:

sorry i didn't reply sooner i had gym

Shalore100:

oh same im in 8th

Shalore100:

not for this one